# Learning Laplacian Eigenvectors: a Pre-training Method for Graph Neural Networks

### v1

@@ A.2 Loss function
Energy loss is order-invariant and rotation invariant (see A.3); for applications in clustering, this is reasonable. However, we would like the model to learn the eigenvectors in their specific order, so we also define absolute energy loss, matching the Rayleigh quotient with the ground-truth eigenvalue:
@@
$\mathcal{L}_{energy\_abs}=\frac{1}{k}\operatorname{Tr}\lvert\hat{U}^{\top}L\hat{U}-\Lambda_{k}\rvert$
@@ A.2.1 Orthogonality
Forced orthogonality, used in [8], imposes orthogonality on the final outputs of the model via QR decomposition. In other words, if $\hat{U}^{\prime}$ is the initial output to the model, $Q$ is an $n\times k$ matrix with orthonormal columns, and $R$ is a $k\times k$ upper triangular matrix, then we achieve the final output $\hat{U}$ as such:
@@
Orthogonality loss, used in [9, 11, 10] imposes a softer constraint, encouraging orthogonality by penalizing the model for producing pairwise similar vectors. This can be written as:
@@
A.3 Energy and eigenvector losses are sign and basis invariant
A.3.1 Definition of basis invariance
@@ A.3.2 Proofs
$\frac{1}{k}\operatorname{Tr}(U^{\top}LU)=\frac{1}{k}\operatorname{Tr}(W^{\top}LW)$
@@
$\frac{1}{k}\operatorname{Tr}(U^{\top}LU)=\frac{1}{k}\operatorname{Tr}(UU^{\top}L)=\frac{1}{k}\operatorname{Tr}(WW^{\top}L)=\frac{1}{k}\operatorname{Tr}(W^{\top}LW)$
A.4 Node feature augmentation
@@
Diffused dirac embeddings encode information about the connectedness of each node and its neighbors. For each node, $m$, we apply the $m^{\text{th}}$ row of the diffusion matrix $P$ to each wavelet $\Psi_{k}$ in our wavelet bank. As above, the difused dirac embedding for node $m$ is the $m^{\text{th}}$ row of the resulting matrix.
@@
Where $U=\begin{bmatrix}\psi_{1}&\dots&\psi_{n}\end{bmatrix}$, with $\psi_{i}$ orthonormal eigenvectors of $L$ and $\Lambda$ is the diagonal matrix of eigenvalues $\lambda_{1},\dots,\lambda_{n}$ of $L$. Any change to the eigenspectrum of $L$, clearly results in a change to $L_{rw}$, and therefore $P$. Since $\Psi_{0}=I-P$, any two graphs with distinct Laplacian eigenspectra will have distinct diffused dirac node embeddings. ∎
A.5 Detailed experimental settings
@@
Method GIN (4.1) GIN (4.1) GIN pre-training (4.2) GPSE (4.3) Pre-training dataset ZINC-12k QM9 (134k) ZINC15 (2M), ChEMBL (456K) MolPCBA (324K) Base architecture GIN GIN GIN MPNN # params 33543 32883 2252210 22075899 # layers of per-node feature update 3 3 2 1 # layers of message passing 4 4 5 20 Hidden dim 60 60 300 512 Activation fn ReLU ReLU ReLU ReLU Dropout 0.1 0.1 0.2 0.2 Batch size 128 128 32 1024 Learning rate 0.001 0.001 0.001 0.005 Optimizer Adam Adam Adam AdamW Scheduler ReduceLROnPlateau ReduceLROnPlateau None CosineWithWarmup patience=5, factor=0.9 patience=20, factor=0.5 - - Pre-Training Epochs 200 100 100 120 Fine Tuning Epochs 500 150 100 - Laplacian norm type Unnormalized Unnormalized Unnormalized Symmetric # eigenvectors predicted 6 6 5 5 Initial features Diffused dirac + Wavelet pos. Diffused dirac + Wavelet pos. Molecule features Random MLP head type(s) Graph-level Graph-level Graph-level, per-node Graph-level, per-node Graph-level MLP max # nodes 40 40 50 50 MLP head # layers 5 5 1 2 MLP head hidden dim 2400 2400 N/A 1600, 32 MLP head activation fn ReLU ReLU N/A ReLU Loss function (and coefficient) 2*Eigenvector + 1*energy 2*Eigenvector + 1*energy 0.25 * Eigenvector + 0.05 * ortho 0.25 * Eigenvector Other features/notes Removed graphs with less than six nodes during pre-training Residual gating, virtual node
A.6 Modifying existing pre-training method
A.6.1 Learning rate tuning
We keep the majority of the settings from [3] the same. For downstream fine-tuning, we tune over 7 learning rates for fair comparison according to [4]. We run each method and learning rate over 3 seeds, and select the learning rate based on mean validation accuracy over all learning rates.
A.6.2 Downstream datasets
We briefly list and cite the five downstream datasets here for reference. The five datasets are the datasets chosen in [4], and are a subset of the eight primary downstream datasets evaluated in [3].
  • BACE: Qualitative binding results [19]
  • BBBP: Blood-brain barrier penetration [20]
  • Tox21: Toxicity data [21]
  • Toxcast: Toxicology measurements [22]
  • SIDER: Database of adverse drug reactions (ADR) [23]
A.7 Modifying existing graph structural encoder
A.7.1 Loss function comparison
We claim that the sum of MAE and cosine similarity loss used in [1] is ill-posed for learning eigenvectors.
To address sign ambiguity, the original GPSE model only learns the absolute value of each Laplacian eigenvector as its target. This improves performance over training without absolute value, but two shortcomings still arise. This (1) provides strictly less information than learning the true eigenvectors, and (2) does not capture basis ambiguity, making it ill-posed for eigenvectors with higher multiplicity (or eigenvectors with small eigengaps).
We demonstrate this claim experimentally in Figure 2. We train the default GPSE settings and architecture using (1) eigenvector loss (ours), and the default cosine similarity + MAE loss (GPSE). We train for 600 epochs on 1% of the MolPCBA dataset and plot eigenvector and energy losses per-epoch. We also plot the loss when outputting random orthogonal vectors for comparison. While training on cosine similarity + MAE achieves better-than-random results, sign and basis ambiguity issues severely limit the model’s ability to learn meaningful eigenvectors.
Figure 2: A comparison of energy loss and eigenvector loss, when the GPSE architecture is trained on (1) eigenvector loss (ours), (2) cosine similarity + MAE [1], and (3) no loss (outputting random vectors). All losses displayed are in natural log scale for visual purposes.

### v2

@@ Appendix E Loss function alternative expressions
Energy loss is order-invariant and rotation invariant (see F); for applications in clustering, this is reasonable. However, we would like the model to learn the eigenvectors in their specific order, so we also define absolute energy loss, matching the Rayleigh quotient with the ground-truth eigenvalue:
@@
$\mathcal{L}_{energy\_abs}=\frac{1}{k}\mbox{\rm Tr}\lvert\hat{U}^{\top}L\hat{U}-\Lambda_{k}\rvert$
@@ E.1 Orthogonality
Forced orthogonality, used in Shaham et al. (2018), imposes orthogonality on the final outputs of the model via QR decomposition. In other words, if $\hat{U}^{\prime}$ is the initial output to the model, $Q$ is an $n\times k$ matrix with orthonormal columns, and $R$ is a $k\times k$ upper triangular matrix, then we achieve the final output $\hat{U}$ as such:
@@
Orthogonality loss, used in Dwivedi et al. (2021); Ma & Zhan (2023); Mishne et al. (2019) imposes a softer constraint, encouraging orthogonality by penalizing the model for producing pairwise similar vectors. This can be written as:
@@
Appendix F Energy and eigenvector losses are sign and basis invariant
F.1 Definition of basis invariance
@@ F.2 Proofs
$\frac{1}{k}\mbox{\rm Tr}(U^{\top}LU)=\frac{1}{k}\mbox{\rm Tr}(W^{\top}LW)$
@@
$\frac{1}{k}\mbox{\rm Tr}(U^{\top}LU)=\frac{1}{k}\mbox{\rm Tr}(UU^{\top}L)=\frac{1}{k}\mbox{\rm Tr}(WW^{\top}L)=\frac{1}{k}\mbox{\rm Tr}(W^{\top}LW)$
$\Box$
Appendix G Node feature augmentation
G.1 Graph Diffusion Operator
@@ G.2 Node Embeddings
Diffused dirac embeddings encode information about the connectedness of each node and its neighbors. For each node, $m$, we apply the $m^{\text{th}}$ row of the diffusion matrix $P$ to each wavelet $\Psi_{k}$ in our wavelet bank. As above, the diffused dirac embedding for node $m$ is the $m^{\text{th}}$ row of the resulting matrix.
@@
These node embeddings are unique up to co-spectrality of the graph Laplacian.
@@
Where $U=\begin{bmatrix}\psi_{1}&\dots&\psi_{n}\end{bmatrix}$, with $\psi_{i}$ orthonormal eigenvectors of $L$ and $\Lambda$ is the diagonal matrix of eigenvalues $\lambda_{1},\dots,\lambda_{n}$ of $L$. Any change to the eigenspectrum of $L$, clearly results in a change to $L_{rw}$, and therefore $P$. Since $\Psi_{0}=I-P$, any two graphs with distinct Laplacian eigenspectra will have distinct diffused dirac node embeddings.    $\Box$
Appendix H Alternative structural pre-training targets
Here, we formally define and provide details for the alternative pre-training targets used in section 4.
  • Node degree: A node-level label representing the degree of each node
  • Local clustering coefficient: A node-level label computing the local clustering coefficient of each node. For a fixed node $u$, the coefficient $C$ is given by:
    $C={\frac{2|\{e_{jk}:v_{j},v_{k}\in N_{u},e_{jk}\in E\}|}{|N_{u}|(|N_{u}|-1)}}.$}
  • RWSE: A node-level label computing self-walk probabilities at varying step counts for the diffusion operator (Dwivedi et al., 2021). In our experiments, we use step counts from the interval $[2,22]$.
  • Cycle counting: A graph-level label computing cycle counts of varying lengths. In our experiments, we count cycles up to length 9.
  • Lap Eigenvalues: A graph-level label computing the $k$-lowest Laplacian eigenvalues $\lambda_{1},\dots,\lambda_{k}$. We use the same $k=6$ as we do with LELM.
For all alternative structural pre-training tasks, we use the same hyperparameters for GIN as displayed in 5, with no initial features and using a standard MAE loss instead of eigenvector + energy loss. We train on the full ZINC dataset. All structural pre-training targets are normalized to have mean 0 and standard deviation 1 across the entire dataset.
Appendix I Enhancing an existing Graph Neural Network pre-training method
I.1 Summary
We augment the existing molecular pre-training methods proposed by Hu et al. (2019) with eigenvector-learning. In particular, Hu et al. (2019) propose node-level pre-training tasks (context prediction and masking) on ZINC15 (Sterling & Irwin, 2015), followed by a graph-level supervised pre-training task on ChEMBL (Mayr et al., 2018; Gaulton et al., 2012). We augment the graph-level supervised pre-training step by adding an additional MLP head to predict eigenvectors, and we evaluate on five downstream datasets based on work by Sun et al. (2022).
Table 4: Test ROC-AUC (%, $\uparrow$) performance on 5 molecular prediction tasks when augmenting an existing pre-training method on a GIN base model. Sup. refers to the original supervised pre-training as implemented by Hu et al. (2019), and Sup.+ refers to supervised training with LELM. Results for no pre-training are taken directly from Sun et al. (2022). All methods are tuned over seven learning rates and averaged over three seeds.
[…]
I.2 Learning rate tuning
We keep the majority of the settings from Hu et al. (2019) the same. For downstream fine-tuning, we tune over 7 learning rates for fair comparison according to Sun et al. (2022). We run each method and learning rate over 3 seeds, and select the learning rate based on mean validation accuracy over all learning rates.
I.3 Downstream datasets
We briefly list and cite the five downstream datasets here for reference. The five datasets are the datasets chosen in Sun et al. (2022), and are a subset of the eight primary downstream datasets evaluated in Hu et al. (2019).
  • BACE: Qualitative binding results Subramanian et al. (2016)
  • BBBP: Blood-brain barrier penetration Martins et al. (2012)
  • Tox21: Toxicity data Mayr et al. (2016)
  • Toxcast: Toxicology measurements Richard et al. (2016)
  • SIDER: Database of adverse drug reactions (ADR) Kuhn et al. (2016)
Appendix J Detailed experimental settings
@@ I.3 Downstream datasets
[…]
For our ablation experiments with node-wise MLP heads, we keep all settings the same, replacing the graph-level MLP, which has an graph-level input dimension of $60\cdot 40=2400$, with a node-level MLP which has an input dimension and hidden dimension of 60.
For ZINC-12k and ZINC, the test/train splits are provided. The QM9 dataset is randomly split into training, validation, and test sets with a ratio of 0.8/0.1/0.1
Appendix K Runtime and compute
All experiments were run using Yale’s Grace cluster. Runs used a single GPU each, with at least 11GB of RAM per GPU. The type of GPU varied between A100, V100, A5000, and RTX5000. For all runs using our main pipeline (comparison against baseline models, comparison against alternative pre-training targets, and ablation on MLP head), pre-training and fine-tuning on all downstream tasks took between 12 to 24 hours on ZINC-full, and 36 to 48 hours on QM9. For experiments enhancing an existing pre-training pipeline, the full pre-training process (including both unsupervised and supervised steps) took around 24 hours, and the downstream fine-tuning across all five datasets for a single seed and model variant took around 15 hours.
Across all results that are included in this paper, we estimate a total of 325 GPU hours spent on the main-body experiments, and an additional 400 GPU hours spent on enhancing an existing GNN pre-training method.
The complete experimental process consumed far more GPU hours, as we initially explored basic models’ abilities to learn Laplacian eigenvectors, and also experimented with a variety of augmentation features before deciding on the two included in the paper. We estimate an additional 800 GPU hours used for initial experimentation and exploration.
Appendix L Visualization of predicted eigenvectors during pre-training
Figure 2: A comparison of predicted eigenvectors (blue) with ground-truth eigenvectors (orange) for a single molecular graph with $n=21$ nodes. These predictions were produced by the standard GIN model on a validation example from the ZINC dataset. Node indices are sorted in increasing order of $\psi_{2}(i)$, and the sign orientation of the predicted vectors $\hat{u}_{i}$ is chosen such that $\hat{u}_{i}\cdot\psi_{i}\geq 0$.
TAG-DS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The claims of the abstract and introduction are that we developed a novel GNN pre-training method that improves performance on structure-based tasks. The empirical results demonstrate the improvements, and the theoretical results justify the design of the framework.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: The paper includes a limitations section, section 5, that discusses the question of whether the framework is suitable for transfer learning.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: All proofs in the paper include the full set of assumptions and all proofs are complete and correct. Proofs are included in sections F and G
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: Code is provided in supplemental material and data is publicly accessible. Moreover, all experimental settings are provided in supplemental material as well for ease of reproducibility.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: Code is provided as part of supplementary material and all data is publicly accessible.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: All experimental details, including data splits, hyperparameters, optimizer, network depth, etc., are provided in the supplemental material in section J.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: Error bars are not reported because it would be too computationally expensive.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We provide a complete overview of compute resources used and time of execution in the appendix in section K.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: Paper does not violate any part of Neurips Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [NA]
  49. Justification: The framework developed in this paper is not tied to any particular application.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [NA]
  54. Justification: All models are trained publicly accessible molecular property prediction datasets that pose no safety risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [NA]
  59. Justification:The paper does not use existing assets.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [NA]
  64. Justification: The paper does not release new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [NA]
  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [NA]
  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [NA]
  79. Justification: LLMs were not used for core methods in this research. The only use of LLMs was for checking grammar in the paper.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.